### v1

@@ Abstract
Multimodal data has become a crucial element in the realm of big data analytics, driving advancements in data exploration, data mining, and empowering artificial intelligence applications. To support high-quality retrieval for these cutting-edge applications, a robust data retrieval platform should meet the requirements for transparent data storage, rich hybrid queries, effective feature representation, and high query efficiency. However, among the existing platforms, traditional schema-on-write systems, multi-model databases, vector databases, and data lakes, which are the primary options for multimodal data retrieval, are difficult to fulfill these requirements simultaneously. Therefore, there is an urgent need to develop a more versatile multimodal data retrieval platform to address these issues.
@@
The study proposes a multimodal feature representation technique that transforms raw multimodal data into representative features, improving effective retrieval of multimodal data.
@@ 1 Introduction
Multimodal data (including video, audio, text, etc.) has reached a critical mass, with 80 percent of worldwide data expected to be in these modalities by 2025 (King, 2020). This variety and sheer volume make "Multimodal Data Retrieval" a core discipline of the more general domain of "Big data" research, which is the cornerstone for supporting big data analysis, driving advancements in data exploration, data mining, and empowering artificial intelligence applications. The efficient handling of large-scale multimodal data is crucial for digging deeper insights and driving innovation across diverse domains, including multimodal knowledge graph (Zhang et al., 2020; Abu-Salih, 2021), personalized recommendation (Wu et al., 2024), disease prediction in healthcare (Andorra et al., 2024; Thukral et al., 2023), and fake news detection in law enforcement (Xue et al., 2021), etc. Moreover, with the evolution and advancement of large-scale models, especially multimodal large-scale language models, AI is endowed with the ability to accept multi-sensory inputs like humans and provide more human-like interactions, which relies heavily on the multimodal data retrieval platform.
The main challenge of a multimodal data retrieval platform is how to efficiently and accurately exploit valuable information from the vast and diverse multimodal data through rich query options along with the concern of storage and management of fast-growing data (Mishra and Misra, 2017). For instance, as shown in Fig 1, we possess multimodal data that includes various types of information such as text, images, video, audio, etc. This data might encompass sports news, dance videos, and life recordings. Our objective is to quickly retrieve the desired information from vast and diverse multimodal data through querying. This requires a retrieval platform that combines efficient storage and querying capability, meeting the following requirements:
Figure 1: Querying in Multimodal Data
(1)Storage layer: Transparent data storage. Multimodal data is typically unstructured and lacks a fixed organization or predefined format, which makes it challenging to store and access data in a seamless manner without requiring users to understand the underlying complexities (Gao et al., 2020). For instance, multimodal data consists of data stored in various formats such as JPG, TXT, MKV, and MP3 files, making the management of such data considerably difficult to hide from the end user.
(2)Query layer: Rich hybrid queries and query-aware mechanism. The effectiveness of information mining for multimodal data largely depends on the richness of the query capabilities provided. Hybrid queries (Wei et al., 2020) refers to search operations over a single vectorized attribute and a set of numerical attributes associated with the searching object, whereas rich hybrid queries extend this concept by supporting more flexible combinations of a broader spectrum of "attribute", which can be vectors or numeric values (see details in Section 4.2). For instance, in the query "Find the information from the soccer match on August 10, 2014, including the images and audio. During the match, the African players in red jerseys repeatedly dribbled through the Asian players in white, earning loud cheers from the audience.", "August 10, 2014" serves as a numeric query, while the subsequent sentence functions as a vector query. Moreover, query-aware refers to the logs of the rich hybrid queries, which involve both execution processes and query results. By maintaining these records, systems can gain insights into query behaviors, which can be used to optimize the effectiveness and efficiency of future multimodal data retrieval operations.
(3) Feature representation layer: Effective feature representation. A common approach to multimodal data retrieval is to "understand" multimodal data by representing it as features and using semantic search to obtain data with corresponding features (Wang et al., 2016). Yet, different multimodal datasets possess diverse and distinct characteristics, and varying query contexts may emphasize different features. Therefore, it is challenging to employ a static process to transform raw data into optimal features and data layouts to meet the requirements of rich hybrid queries. Instead, a query-aware process is necessary to obtain the optimal features.
(4)Index layer: High query efficiency. Different users and application scenarios usually have diverse query patterns. Journalists, wish to quickly retrieve sports news from the multimodal data, whereas for bloggers, their queries are more likely related to life recordings. It has emerged as a significant challenge to devise and update system’s index structure to adapt to different query workloads and efficiently manage computational costs.
@@
In response to the growing need for inferring meaningful insights from large amounts of multimodal data, data platforms have been evolving to meet the demands of multimodal data retrieval. Existing retrieval platforms can be categorized into traditional schema-on-write systems, multi-model databases, vector databases, and data lakes. (1) Traditional schema-on-write (i.e., store data in predefined schema) systems, like relational databases, are inflexible for handling multimodal data transparent storage and hybrid queries due to their fixed schema constraints when managing data. (2) Multi-model databases have been proposed to integrate several schema models within a single, integrated backend, allowing them to handle several data formats within one unified system (Lu and Holubová, 2019; Lei et al., 2024). While these databases can offer transparent storage for certain data modalities (e.g., text), their storage capabilities are often limited when dealing with a comprehensive range of multimodal data. Additionally, their functionality and efficiency are constrained by the lack of rich query options and robust indexing structure. (3) Vector databases are designed specifically for efficient retrieval of high-dimensional vector data through vector similarity searches or simple hybrid queries (Pan et al., 2024), often overlooking the storage of the original multimodal data and the process of feature representation. Despite their specialized design, vector databases face significant challenges in maintaining efficiency when managing vectors with higher dimensionality (Taipalus, 2024). (4) Data lakes are repositories with data from disparate sources that are transparently stored in their original, raw format (Khine and Wang, 2018; Hai et al., 2023). Organizations often regard data lakes as a cost-effective solution for storing vast amounts of data and enabling additional computational functionalities to support retrieval operations, but most of them lack support for rich hybrid queries (Ren et al., 2021b, a; Xiao et al., 2022). Furthermore, the efficiency of queries in data lakes is often compromised due to the insufficient focus on the feature representation and indexing processes of raw data. Therefore, in this study, we explore how to leverage the potential of the data lake to propose a more full-fledged multimodal data retrieval platform based on data lake. In addition to providing infrastructural transparent storage and rich hybrid queries, we mainly focus on improving the effectiveness and efficiency of the data retrieval platform.
@@
For improving effectiveness, the robustness of the feature representation is critical. This involves transforming raw data into representative features and optimal data layout, in order to benefit subsequent tasks such as indexing and querying. Various embedding approaches have been proposed, ranging from single-modality embedding methods like doc2vec (Le and Mikolov, 2014) and graph2vec (Grohe, 2020) to advanced multi-modality embedding techniques such as CLIP (Radford et al., 2021). Despite the availability of numerous embedding methods, it remains a challenging task to measure and enhance these features to make them semantically rich and computationally efficient for retrieval tasks. Existing feature measurement and enhancement methods primarily focus on optimizing features for building machine learning models. Feature measurement techniques aim to identify features best suited for model construction. Feature enhancement methods, such as imputation (Awawdeh et al., 2022), discretization (Dougherty et al., 1995), and feature creation (Rohrer, 2011; Falcon et al., 2022), target at improving model training performance. These methods are not tailored to the scenario of multimodal data retrieval where the goal is enabling the multimodal retrieval platform to perform queries with greater accuracy and speed.
@@
For improving efficiency, the quality of indexes is crucial for achieving accurate and fast data retrieval. (1) Traditional multi-dimensional indexes such as grid file (Nievergelt et al., 1984) and R-tree (Guttman, 1984) are designed to efficiently support numeric queries, particularly on low-dimensional data, while having difficulties when faced with high-dimensional data. (2) Vector similarity indexes such as ANNOY (Bernhardsson, 2015) and HNSW (Malkov and Yashunin, 2018), excel in the realm of high-dimensional vector data indexing through distance computation in vector space to find similar vectors. Nevertheless, they cannot handle hybrid queries that combine numeric queries and vector queries, and increasing data dimensionality and volume can significantly reduce their query efficiency. (3) High-dimensional learned indexes extends traditional multi-dimensional structures by employing machine learning models to quickly and accurately find expected results (Al-Mamun et al., 2024). This method offers a pathway for indexes to optimize their internal structures by learning data characteristics and query behaviors, especially in scenarios involving large datasets. But still, current works in learned indexes primarily focus on proving efficiently low-dimensional numeric queries. So far, exploring an efficient query method that supports both vector and numeric queries, as well as their combinations, remains a challenge yet to be fully resolved.
@@
Therefore, we summarize the key problems to be solved in this paper as:
@@
    How to achieve transparent and seamless storage, as well as rich hybrid queries, for multimodal data retrieval?
@@
    How to convert multimodal data into an optimal feature representation to enhance query result quality?
@@
    How to support efficient multimodal data retrieval?
To address these problems, we present a multimodal data retrieval platform with query-aware feature representation and learned index based on data lake (MQRLD), which supports transparent data storage, and rich hybrid queries empowered by our high-dimensional learned index that benefits from effective multimodal data feature representation and a query-aware mechanism to deliver optimal performance. The main contributions are as follows:
@@
    We provide a seamless management and flexible access method for multimodal data, including the utilization of data lake technology to support multimodal data storage, and the usage of multimodal open API to support rich hybrid queries, based on which an efficient query-aware mechanism is built to enhance query performance.
@@
    We propose a multimodal data feature representation technique designed to transform multimodal raw data into representative features and optimal data layout by perceiving both data characteristics and query behaviors, in order to facilitate index building, thereby enhancing the query effectiveness.
@@
    We present a high-dimensional learned index that can adaptively optimize its inner structure under different feature data layouts and various task scenarios to support flexible and efficient data retrieval of multimodal data.
@@ 2 Related Work
In view of the requirements mentioned above, we summarize the current works on data retrieval platforms, feature representation techniques, and indexes. We first enumerate platforms supporting multimodal data retrieval in Section 2.1, focusing on architecture and functionality which include storage, query options, feature representation functionality, and index type. Then in Section 2.2, we introduce the existing techniques in feature representation domain. Finally, we conduct a more detailed comparison of high-dimensional indexes in Section 2.3, focusing on functionality and performance, which include index structure, query types, and whether data-aware (considering multimodal data characteristics) or query-aware (considering query behaviors) techniques are applied to improve index efficiency.
@@ 2.1 Related work on data retrieval platform
Table 1: Data Retrieval Platforms.
@@
PostgreSQL (Obe and Hsu, 2017) schema-on-write system traditional multi-dimensional index/ vector similarity index
@@
(Belgundi et al., 2023) multi-model database traditional multi-dimensional index
Azure Cosmos DB (Guay Paz, 2018) multi-model database traditional multi-dimensional index
OrientDB (Ritter et al., 2021) multi-model database traditional multi-dimensional index
@@
HMDFF (Ren et al., 2021b) data lake traditional multi-dimensional index
@@
Traditional schema-on-write systems require a fixed schema for data storage, organize numerical data in columns, and utilize traditional one-dimensional or multi-dimensional indexes for data retrieval. This structure is well-suited for managing numerical data, for example, PostgreSQL performs effectively in supporting bank transactions (Roskladka et al., 2019). However, this fixed schema approach is less adaptable to high-dimensional data. Although plugins like pgvector (pgvector, 2024) have been introduced to handle vector-level data processing and similarity search, they are unable to deliver high performance and scalability required for large-scale applications.
@@
Standard vector databases, such as Faiss (Johnson et al., 2019) and Pinecone (Pinecone, 2024), focus solely on the storage and retrieval of vector-format data and use only vector similarity indexes. This limitation prevents them from executing hybrid queries. Extensions like ADAMpro (Giangreco and Schuldt, 2016), Cottontail DB (Gasser et al., 2020), and Milvus (Wang et al., 2021) offer additional functionalities, such as hybrid queries and feature embedding. However, they exhibit significant flaws: Firstly, they lack explicit processing from raw data to vectors and neglect the storage of raw data. Secondly, despite providing basic feature embedding functionality, they cannot guarantee the quality of features, thus impacting search effectiveness. Thirdly, although they can provide simple hybrid queries, this comes at the expense of significantly reduced query performance. This heavy query process leads to exponential degradation in performance as the vector dimension increases, making it difficult to handle more complex queries.
@@
Existing retrieval platforms perform well within their specific application areas; however, none of them are likely to solve all the challenges related to multimodal data retrieval. Moreover, none of these platforms improve query performance using query-aware or learned index techniques, resulting in suboptimal query efficiency.

### v2

@@ Abstract
Multimodal data has become a crucial element in the realm of big data analytics, driving advancements in data exploration, data mining, and empowering artificial intelligence applications. To support high-quality retrieval for these cutting-edge applications, a robust multimodal data retrieval platform should meet the challenges of transparent data storage, rich hybrid queries, effective feature representation, and high query efficiency. However, among the existing platforms, traditional schema-on-write systems, multi-model databases, vector databases, and data lakes, which are the primary options for multimodal data retrieval, make it difficult to fulfill these challenges simultaneously. Therefore, there is an urgent need to develop a more versatile multimodal data retrieval platform to address these issues.
@@
The study proposes a multimodal feature representation technique that converts raw multimodal data into representative features and optimal data layouts, improving the effective retrieval of multimodal data.
@@ 1 Introduction
Multimodal data, including both structured and unstructured data, continuously influxes in vast volumes from various sources. These data often need to be integrated to extract comprehensive and meaningful insights (Lymperaiou and Stamou, 2024). This makes "Multimodal Data Retrieval" a core discipline of the more general domain of "Big data" research. For instance, in power systems, various devices continuously collect multimodal data (supervisory control and data acquisition logs, meteorological data, satellite remote sensing images, surveillance videos, etc) from diverse sources, and the growing interactive demand response requires power companies to efficiently extract target data from massive datasets (Zhao et al., 2022). Similarly, large e-commerce platforms like Taobao and Amazon can provide millions of users with billions of items (the products’ names, prices, images, promotional videos, etc), and the users usually search for their desired products on these platforms (Belem et al., 2020; Zheng et al., 2021; Rasappan et al., 2024). The efficient handling of large-scale multimodal data is crucial for digging deeper insights, such as big data analysis, data exploration and data mining, and driving innovation across diverse domains, including recommendation (Wu et al., 2024), disease prediction in healthcare (Andorra et al., 2024; Thukral et al., 2023), fake news detection in law enforcement (Xue et al., 2021), etc. Moreover, with the evolution and advancement of large-scale models, especially multimodal large-scale language models, AI is endowed with the ability to accept multi-sensory inputs like humans and provide more human-like interactions.
All these applications heavily rely on multimodal data retrieval, as it can significantly impact their real-world performance. In multimodal data retrieval, the data can be handled as multimodal objects (MMOs), which consist of complex information that combines structured data (usually represented as numerical attributes) and unstructured data (usually represented as vector features). As the real-world scenario shown in Fig 1, multimodal data retrieval platforms are required to support complex joint queries (we call them rich hybrid queries), with the query results returned in the form of MMOs. However, existing multimodal data retrieval platforms, such as traditional schema-on-write systems, multi-model databases, vector databases, and data lakes, exhibit limited capabilities in achieving transparent storage and supporting rich hybrid queries simultaneously. Moreover, their query efficiency and accuracy tend to degrade when handling large-scale datasets. Therefore, the main objective of research on multimodal data retrieval platforms is to develop a platform that can efficiently and effectively retrieve MMOs within large-scale multimodal data through rich hybrid queries along with the concern of transparent data storage (Mishra and Misra, 2017). To achieve this, retrieval platforms must overcome the following challenges:
Figure 1: Querying in multimodal data, including structured query attributes ("10\$-20\$" and "0-24 hours") and vector query attributes ("CupColor.jpg" and "CupDescription.mp3").
(1)Storage layer: Transparent data storage. Users usually aim to retrieve target MMOs through query statements. As shown in Fig 1, the user is more concerned with obtaining all the information of his desired products (MMOs), which includes the products’ names, prices, images, promotional videos, etc. However, multimodal data often involves heterogeneous sources, varied modalities such as text, images, and videos, and complex interrelationships among different types of information, lacking a fixed organization or predefined format. This makes it challenging to store and access these data in a seamless manner without requiring users to understand the underlying complexities (Gao et al., 2020). To address this limitation, a multimodal data retrieval platform should enable the conversion and tracking of information within MMOs, thereby supporting transparent data storage.
(2)Query layer: Rich hybrid queries and query-aware mechanism. The richness of multimodal data exploration and mining depends largely on the query functionality, while the retrieval performance relies on the ability of the platform to perceive query behaviors. Compared to the hybrid queries proposed in existing research (Wei et al., 2020), which support queries containing a single vector and a set of structured attributes, many real-world scenarios require more complex joint queries. These queries are referred to as rich hybrid queries, which support more flexible combinations of a broader spectrum of "attributes", whether they can be structured or vector attributes (as detailed in Section 4.2). For instance, in Fig 1 rich hybrid query statement, "priced between 10 and 20" and "delivered tomorrow" serve as structured query attributes, while the "CupColor.jpg" and "CupDescription.mp3" as vector query attributes. Additionally, real-world query behavior is often skewed rather than average. To enhance the performance of multimodal data retrieval, platforms must also be aware of these query behaviors. Therefore, a multimodal data retrieval platform should support rich hybrid queries and be capable of perceiving query behaviors to provide effective and efficient responses.
(3) Feature representation layer: Effective feature representation. A common approach to multimodal data retrieval is to "perceive" and "understand" multimodal data by representing it as features and obtaining data with corresponding features (Wang et al., 2016). Given the vast amount of multimodal data, where different types emphasize distinct information, flexible feature embedding methods and unified measurement techniques are essential to accommodate varying datasets. Additionally, query behaviors often prioritize specific features, impacting retrieval precision and recall. This requires a data-aware and query-aware feature enhancement approach to dynamically adjust data layouts. Therefore, a multimodal data retrieval platform should have a unified framework for feature embedding, measurement, and enhancement to achieve effective feature representation.
(4)Index layer: High query efficiency. Efficient multi-dimensional and high-dimensional indexes are a key solution for supporting efficient rich hybrid queries on MMOs. Therefore, a multimodal data retrieval platform needs an indexing structure capable of handling rich hybrid queries that include both vector features and structured attributes. Moreover, different users usually have diverse query patterns. For instance, on an e-commerce platform, new parents typically search for baby products, whereas for students, their queries are typically focused on study supplies. Consequently, designing a high-dimensional indexing structure tailored for rich hybrid queries, along with mechanisms to adaptively update the index structure based on query workloads to reduce computational costs, has become a significant challenge.
@@
For fulfilling the storage and retrieval functions for multimodal data, data platforms have been evolving to meet the growing need for inferring meaningful insights from these data. Existing retrieval platforms can be categorized into traditional schema-on-write systems, multi-model databases, vector databases, and data lakes. (1) Traditional schema-on-write systems such as PostgreSQL (Obe and Hsu, 2017) and MySQL (Oracle Corporation, 2023), store data in predefined schemas and are commonly used for structured data management and transactional operations. These systems are often inflexible when handling multimodal data, as they are constrained by their fixed schema, making it difficult to realize transparent storage and support complex hybrid queries. (2) Multi-model databases have been proposed to integrate several schema models within a single, integrated backend, allowing them to handle several data formats within one unified system (Lu and Holubová, 2019; Lei et al., 2024). For example, MongoDB (MongoDB, Inc., 2023) can offer transparent storage for certain data modalities, but its storage capabilities are limited when dealing with wilder range of multimodal data, such as video and audio. Additionally, their functionality and efficiency are constrained when handling rich hybrid queries, especially for high-dimensional data. (3) Vector databases are designed specifically for efficient retrieval of high-dimensional vector data through vector similarity searches(Pan et al., 2024). These platforms, such as Faiss (Johnson et al., 2019), are specialized in managing vector data, which is often used for machine learning or AI-powered applications. However, they typically focus only on vector queries and neglect the original multimodal data storage and feature representation process. While highly effective for tasks involving a single type of data (e.g., image retrieval or document search), vector databases may struggle with rich hybrid queries that involve multiple data attributes, such as a combination of vectors and structured attributes (Taipalus, 2024). (4) Data lakes, such as Apache Hudi (Hudi, 2021), are large repositories that transparently store data in its original, raw format sourced from diverse origins(Khine and Wang, 2018; Hai et al., 2023). Data lakes are commonly seen as a cost-effective solution for storing vast amounts of raw data, and support adding additional computational layers, but most of them lack support for rich hybrid queries(Ren et al., 2021b, a; Xiao et al., 2022). Furthermore, the efficiency of queries in data lakes can be compromised due to insufficient attention to feature representation and indexing of raw data. Therefore, in this study, we explore how to leverage the potential of the data lake to propose a more full-fledged multimodal data retrieval platform. In addition to providing infrastructural transparent storage and rich hybrid queries, we focus on improving the effectiveness and efficiency of the data retrieval platform.
@@
For improving the effectiveness of multimodal data retrieval, feature representation is critical. Effective feature representation involves converting raw data into representative features and optimal data layout. Therefore, a retrieval platform must convert data into optimal representative features through feature embedding and measurement, followed by feature enhancement to optimize the data layout. However, most existing multimodal data retrieval platforms rely on fixed methods for feature embedding and layout adjustments, failing to integrate feature embedding, measurement, and enhancement into a unified framework for achieving effective feature representation. Specifically, some progresses have been made in feature embedding, measurement, and enhancement separately. Current research has proposed various embedding methods, ranging from single-modality embedding methods like doc2vec (Le and Mikolov, 2014) and graph2vec (Grohe, 2020) to advanced multi-modality embedding techniques such as CLIP (Radford et al., 2021). Existing feature measurement and enhancement methods primarily focus on optimizing features for building machine learning models, whereas feature measurement techniques aim to identify features best suited for model construction, and feature enhancement methods, such as imputation (Awawdeh et al., 2022), discretization (Dougherty et al., 1995), and feature creation (Rohrer, 2011; Falcon et al., 2022), target at improving model training performance. Although these methods perform well on specific scenarios, they are not well-suited for enabling more effective retrieval on multimodal data retrieval platforms.
@@
To improve the efficiency of multimodal data retrieval, the index must be query-aware and support rich hybrid queries. Current indexing methods mainly include traditional multi-dimensional indexes, vector similarity indexes, and multi-dimensional learned indexes. (1) Traditional multi-dimensional indexes such as grid file (Nievergelt et al., 1984) and R-tree (Guttman, 1984) are designed to efficiently support numeric queries, particularly on low-dimensional data, while having difficulties when faced with rich hybrid queries and lacked query-aware mechanism. (2) Vector similarity indexes such as ANNOY (Bernhardsson, 2015) and HNSW (Malkov and Yashunin, 2018), excel in the realm of high-dimensional vector data indexing through distance computation to find similar vectors. Nevertheless, they cannot handle rich hybrid queries performed on MMOs, and increasing data records can significantly decrease their query efficiency. (3) Multi-dimensional learned indexes extends traditional multi-dimensional structures by employing machine learning models to quickly find expected results (Al-Mamun et al., 2024). This method offers a query-aware mechanism for indexes to optimize their internal structures by learning data layout and query behaviors. But still, current works in learned indexes primarily focus on efficiently proving low-dimensional queries, failing to achieve rich hybrid queries. Current data retrieval platforms typically achieve rich hybrid queries by constructing two separate indexes: a high-dimensional vector index for unstructured data and a multi-dimensional index for structured data (Wei et al., 2020). So far, exploring an efficient query method that supports rich hybrid queries with a query-aware mechanism, remains a challenge yet to be fully resolved.
@@
Therefore, we summarize the key problems of multimodal data retrieval platforms to be solved in this paper as:
@@
    How to achieve transparent data storage and rich hybrid queries, where user queries involve flexible combinations of structured and unstructured attributes, and the final results are returned as MMOs?
@@
    How to implement a unified feature representation technique for feature embedding, measurement, and enhancement to convert multimodal data into optimal features and data layouts, thereby improving the effectiveness of query results?
@@
    How to implement a query-aware high-dimensional index that manages massive vectors and structured data, supporting efficient multimodal data retrieval?
To address these problems, we present a multimodal data retrieval platform with query-aware feature representation and learned index based on data lake (MQRLD), which supports transparent data storage and rich hybrid queries empowered by our high-dimensional learned index that benefits from effective multimodal data feature representation and a query-aware mechanism to deliver optimal performance. The main contributions are as follows:
@@
    We provide a seamless and flexible access method for multimodal data, including the utilization of data lake technology to support transparent multimodal data storage, and the usage of multimodal open API to support rich hybrid queries which return query results in the form of MMOs, based on which a query-aware mechanism is built to enhance query performance. (Section 4)
@@
    We propose a unified multimodal data feature representation technique, which is designed to convert multimodal raw data into representative features and optimal data layout by perceiving and understanding both data characteristics and query behaviors, in order to facilitate index building, thereby enhancing the query effectiveness. (Section 5)
@@
    We present an efficient high-dimensional learned index built on feature representation that can adaptively optimize its inner structure under different feature data layouts and various task scenarios to support flexible and efficient data retrieval of multimodal data. (Section 6)
  4. 4.
    We conduct extensive performance evaluations using real and synthetic datasets. The results demonstrate that MQRLD outperforms other methods in multimodal data retrieval tasks, showing its ability to significantly enhance retrieval performance in complex environments. (Section 7)
@@ 2 Related Work
In view of the challenges mentioned above, we summarize the current works on data retrieval platforms, feature representation techniques, and typical multi-dimensional and high-dimensional indexes used in data retrieval platforms. We first enumerate data retrieval platforms in Section 2.1, focusing on architecture and functionality which include storage, query options, feature representation functionality, and index type. Then in Section 2.2, we introduce the existing techniques in feature representation domain. Finally, we conduct a more detailed comparison of multi-dimensional and high-dimensional indexes in Section 2.3, focusing on Structure and Functionality, which include index structure, query types, and whether data-aware or query-aware techniques are applied to improve index efficiency.
@@ 2.1 Related Work on Data Retrieval Platform
Table 1: Comparison of data retrieval platforms on Type, Storage Layer, Query Layer, Feature Representation Layer, and Index Layer.
@@
PostgreSQL (Obe and Hsu, 2017) schema-on-write system multi-dimensional index/ vector similarity index
@@
(Belgundi et al., 2023) multi-model database multi-dimensional index
Azure Cosmos DB (Guay Paz, 2018) multi-model database multi-dimensional index
OrientDB (Ritter et al., 2021) multi-model database multi-dimensional index
MongoDB (MongoDB, Inc., 2023) multi-model database multi-dimensional index
@@
HMDFF (Ren et al., 2021b) data lake multi-dimensional index
@@
Traditional schema-on-write systems require a fixed schema for data storage, organize data in columns, and utilize traditional one-dimensional or multi-dimensional indexes for data retrieval. This structure is well-suited for managing structured data, for example, PostgreSQL performs effectively in supporting bank transactions (Roskladka et al., 2019). However, this fixed schema approach is less adaptable to high-dimensional data. Although plugins like pgvector (pgvector, 2024) have been introduced to handle high-dimensional vector data processing and similarity search, they are unable to deliver the high performance and scalability required for large-scale applications.
@@
Standard vector databases, such as Faiss (Johnson et al., 2019) and Pinecone (Pinecone, 2024), focus solely on the storage and retrieval of vector-format data and use only vector similarity indexes. This limitation prevents them from executing rich hybrid queries. Extensions like ADAMpro (Giangreco and Schuldt, 2016), Cottontail DB (Gasser et al., 2020), and Milvus (Wang et al., 2021) offer additional functionalities, such as simple hybrid queries(e.g. a structured attribute and a vector feature) and feature embedding. However, they exhibit significant flaws: 1. Firstly, they neglect the storage of raw data and need to artificially distinguish between the processing of different data types. 2. Secondly, they lack MMO management processing, making it explicit to track back original data. 3. Thirdly, despite providing basic feature embedding functionality, they cannot guarantee the quality of features, thus impacting search effectiveness. 4. Finally, although they can support simple hybrid queries, this comes at the expense of significantly reduced query performance. This heavy query process leads to exponential degradation in performance as the data records increase, making it difficult to handle large-scale data.
@@
Existing data retrieval platforms perform well within their specific application areas; however, none of them are likely to solve all the challenges related to multimodal data retrieval. Moreover, none of these platforms improve query performance using query-aware or learned index techniques, resulting in suboptimal query efficiency.